OpenStudy (anonymous):

Identify the vertex for the graph of y = −2x2 − 12x + 5.

OpenStudy (anonymous):

vertex: y = −2x^2 − 12x + 5 y = −4x − 12 set it to zero 0=-4x-12 12=-4x -3=x

OpenStudy (anonymous):

You vertex will be (-3,y) plug -3 to the original equation and you got yourself a y

OpenStudy (anonymous):

need more help?

OpenStudy (anonymous):

I take that as a yes ok then when x=-3 y=23 vertex is then (-3,23)

OpenStudy (anonymous):

THANK YOU SO MUCH!!! :)